1. In testing for the effectiveness of an antacid compound, 5.0 g hydrochloric acid is mixed with magnesium hydroxide.
  2. Write a balanced equation for the reaction.                       

a.Determine the number of moles of magnesium hydroxide required to react with the acid.                                               

b.If there is 24.0 g magnesium hydroxide in the lab, determine if it will be enough to react with all the acid.       


1
Expert's answer
2021-04-25T03:50:13-0400

"Mg(OH)_2(s) + 2HCl(aq) \u2192 2H_2O(l) + MgCl_2(aq)"


Moles of HCl"=\frac{5.0}{36.458}=0.137mol"


Moles of "Mg(OH)_2=\frac{0.137}{2}=0.0685mol"



Moles would be"= \frac{24}{58}=0.414mol"


Yes it would be enough